OpenStudy (clamin):

MEDAL!! PLEASE HELP!! 6x-14x+16=-35+11

OpenStudy (dtan5457):

Solving for x?

OpenStudy (clamin):

yes

OpenStudy (dtan5457):

so just isolate x and put it to one side

OpenStudy (dtan5457):

6x-14x+16=-35+11 6x-14x=-35+11-16 simplify that

OpenStudy (clamin):

yea i know how to solve it but i got diff. answer

OpenStudy (dtan5457):

what's the answer choices and what did you get?

OpenStudy (clamin):

ohh im sorry i didnt see the 16..my bad

OpenStudy (clamin):

5 is the answer

OpenStudy (dtan5457):

yes, it is.